An arrow is shot from a bow at an angle of 25 ° above the horizontal with an initial speed of 45 m / s. Find the horizontal and vertical components of the arrow's initial velocity.
1
Expert's answer
2020-12-04T09:36:01-0500

Let's find the horizontal and vertical components of the arrow's initial velocity:


"v_{0x}=v_0cos\theta=45\ \dfrac{m}{s}\cdot cos25^{\circ}=40.78\ \dfrac{m}{s},""v_{0y}=v_0sin\theta=45\ \dfrac{m}{s}\cdot sin25^{\circ}=19.01\ \dfrac{m}{s}."

Answer:

"v_{0x}=40.78\ \dfrac{m}{s}, v_{0y}=19.01\ \dfrac{m}{s}."